# PerfTracker: Online Performance Diagnosis for Large Model Training

### v3

@@ 6.3. System Overhead
Overhead in different model configurations. Since profiling and data generation are the two components executed in the main process of LMT (Figure 16), Table 3 evaluates them thoroughly in different configurations. In general, the time spent on data generation correlates positively with the number of function executions on a single worker within the profiling window. Increasing the degree of parallelism (e.g., the tensor parallelism, TP) results in more fragmented model training. This fragmentation leads to more function execution events, increasing the data generation time.
Table 3. Overhead in different model configurations
@@
model size tp pp training profiling generate data
@@
In most configurations, profiling does not introduce overhead to training. However, when the model is small and the parallelism parameters are large (e.g., GPT-3 7B with TP=2; GPT-3 13B with TP=4 or 8), there is a major increase in training time. This is because such configurations impose a higher load on the CPU; as profiling requires CPU computation, it causes contention. Fortunately, these impractical configurations are never used in production. Note: PerfTracker is still effective in these configurations and the overhead only happens during the profiling window.
6.4. Comparisons with State of the Arts
We compared PerfTracker with state-of-the-art techniques described in recent papers (Jiang et al., 2024; Deng et al., 2024; Dong et al., 2024), as well as existing profiling tools. Table 4 shows the comparison.
For online monitors, we compare PerfTracker with techniques in MegaScale (Jiang et al., 2024) that records all CUDA events in a timeline (to reveal slow GPU computations) and millisecond- to second-level monitoring of RDMA. Such techniques only trigger alerts or display statistical results; root cause localization (especially for network issues) is usually completed manually. As such techniques do not collect Python function events, they cannot address code-level issues. Similarly, the technique in Minder (Deng et al., 2024) only collects hardware metrics but is unaware of LMT executions. It focuses on LMT crashes, not performance issues. The technique in C4 (Dong et al., 2024) collects runtime information from the communication library, and hence only addresses communication problems.
We also compare PerfTracker with directly applying profilers like Nsight Systems (nsy, 2024) and Torch Profiler (tor, 2024). Nsight provides a Post-Collection Analysis API, which can take profiling data as input for custom analysis. Torch Profiler can dump profiling data as JSON files for offline analysis. Not surprisingly, directly using these profilers online incurs significant overhead in production LMT—even reading fine-grained profiling data is untenable.
Table 4. Comparisons with state-of-the-art approaches
Technique Problems Diagnostic time for 10,000-GPU LMT
C1P1 C1P2 C2P1 C2P2 C2P3
MegaScale (Jiang et al., 2024) online
Minder (Deng et al., 2024) online
C4 (Dong et al., 2024) online
Nsight System ¿1.5 days (offline)
Torch Profiler ¿3.5 days (offline)
PerfTracker 3 min (online)
Table 4 shows the comparison of PerfTracker and the reference techniques in terms of the troubleshooting ability and diagnosis time using the problems in §6.1 and §6.2 (e.g., C1P1 represents Problem 1 of Case 1 in §6.1). Since existing online monitors have incomplete data sources, they cannot troubleshoot many problems. While profilers (Nsight and Torch Profiler) can cover more, loading profiling data of all workers requires days. Since they do not provide troubleshooting logic, we only count the time for data loading in Table 4. Together with troubleshooting, they required even more time.
7. Limitations and Discussions
Scope. PerfTracker does not ensure automatically identifying the root causes of all possible performance issues. (1) While for most performance issues, PerfTracker’s output of the problematic functions and their corresponding runtime behavior patterns is sufficient to determine or closely approximate the root cause, some Python functions and CPU kernels have complex logic. Therefore, even when PerfTracker identifies that these functions have problems (e.g., excessively long execution times), analyzing the root causes still requires manual inspection of their code. (2) As described in §4.3, PerfTracker does not consider any runtime behavior patterns of GPU computation kernels as unexpected. Consequently, if a user manually implements an inefficient GPU computation kernel that runs slowly on all workers, it will not be detected by the distance from expectation or differential distance described in §4.3. PerfTracker requires prior knowledge (e.g., an expected range for this kernel function) to diagnose the issue.
LMTs with non-NVIDIA GPUs or non-PyTroch frameworks. In the clusters where PerfTracker has been deployed, all LMTs (e.g., in Table 2) use Nvidia GPUs. Additionally, PerfTracker utilizes nsys to sample hardware metrics, which is only applicable to Nvidia GPUs. In fact, PerfTracker has also been tested on clusters with non-Nvidia GPUs. In these environments, PerfTracker can still use Torch Profiler to collect Python functions, CPU operations, memory operations, and CUDA kernels, but hardware metrics must be collected using tools specific to other GPU models. Therefore, if the sampling tools provided by these GPUs are less precise than nsys, it may affect the ability to diagnose some hardware-related performance issues. The ability to diagnose code-level problems remains unchanged. In addition, PerfTracker only supports PyTorch at present since PyTorch is the dominant framework. However, the insight and the design of this paper are generic to any frameworks (like TensorFlow (Abadi et al., 2016)).
8. Related Works
Hardware monitoring tools. GPUs can be monitored by a series of tools developed by NVIDIA. NVML (nvm, 2023) is a comprehensive, C-based API library. It provides a standardized interface for users to interact with the GPUs, allowing them to actively sample the status of these devices. DCGM (dcg, 2024) is specifically designed for monitoring GPUs in large data centers. It offers a centralized and scalable monitoring method, making it highly suitable for enterprise environments with extensive GPU deployments. CPUs can be monitored by PCM (pcm, 2024), an application programming interface for monitoring performance and energy metrics of Intel processors. For monitoring NICs (e.g., Mellanox), mstflint (mst, 2024) is developed to access packet or byte-level throughput data.
LMT profiling and analysis. Nsight Systems and Nsight Compute provide comprehensive performance analysis and profiling capabilities. Nsight Systems offers a holistic view of CPU, GPU, and memory usage, while Nsight Compute focuses on detailed kernel performance metrics for CUDA applications, helping developers identify and optimize computational bottlenecks. Torch Profiler is a tool within the PyTorch ecosystem that assists in optimizing PyTorch models by providing detailed insights for each function.
Fast recovery. Fast recovery for LMT is currently achieved through periodic (Microsoft, 2023; T. L. et al., 2023) or incremental (Eisenman et al., 2022) checkpointing of the LMT state to the host memory/disks. Checkfreq (Mohan et al., 2021) mitigates the checkpointing overhead by overlapping computation with the checkpoint process, and employs an adaptive runtime checkpointing frequency through profiling. Gemini (Wang et al., 2023) transfers checkpointing data to multiple hosts to improve robustness. It performs the transmission during the gaps between collective communications. Tanmaey et al. (Gupta et al., 2024) only make a checkpoint after an error occurs, and the faulty rank is recovered from the state of the healthy ranks a minibatch before. This avoids the need to re-initialize the entire training job state. There are other approaches to reduce the checkpointing overhead by asynchronous checkpointing (Nicolae et al., 2020) and partial checkpointing (Maeng et al., 2021).
9. Conclusion
This paper presents PerfTracker, an online diagnosis system for LMT. PerfTracker presents the process of performance degradation detection, behavior pattern summarization, and performance issue localization, which utilize runtime behavior patterns of functions to make diagnosis. We deploy PerfTracker in real-world production GPU clusters and solve sophisticated performance issues for real users.
APPENDIX
Appendix A Examples of LMT executions
Figure 18. The runtime behavior of an MoE training.
Figure 19. The python function behaviors of a forward step in the MoE training.
Figure 20. The kernel function (includes both computation and communication) behaviors of a forward step in the MoE training.
@@ Appendix A Examples of LMT executions
Figure 19 and 20 zoom into one of the forward phase in Figure 18. It shows that in the forward phase, low-level functions are also executed repeatedly, and their behaviors (e.g., duration of execution) are highly similar across multiple executions.
Notice that in some configurations, the function runtime behaviors may be slightly different between workers. For example, in pipeline parallel, the workers in the middle pipeline stages require extra send_recvs compared with the first and the last stage. However, function behaviors of workers in the same pipeline stage (in different data parallel groups for example) are still expected to be identical. As a result, the idea of making comparisons of runtime behavior patterns for performance diagnosis still makes sense in these cases.

### v4

@@ Appendix D Profiling Overhead in Different Model Configurations
As explained in §6.4, with Eroica, profiling and data generation are the two components executed in the main process of LMT (Figure 16), Table 4 evaluates them systematically in different LMT configurations. In general, the time spent on data generation correlates positively with the number of function executions on a single worker within the profiling window. Increasing the degree of parallelism (e.g., the tensor parallelism, TP) results in more fragmented model training. This fragmentation leads to more function execution events, increasing the data generation time.
Table 4: Overhead in different model configurations
@@
model size tp $p$ training profiling generate data
@@
In most configurations, profiling does not introduce overhead to training. However, when the model is small and the parallelism parameters are large (e.g., GPT-3 7B with TP=2; GPT-3 13B with TP=4 or 8), there is a major increase in training time. This is because such configurations impose a higher load on the CPU; as profiling requires CPU computation, it causes contention. Fortunately, these impractical configurations are never used in production. Note: Eroica is still effective in these configurations and the overhead only happens during the profiling window.
Appendix E Examples of LMT executions
@@
Figure 22 and 23 zoom into one of the forward phase in Figure 21. It shows that in the forward phase, low-level functions are also executed repeatedly, and their behaviors (e.g., duration of execution) are highly similar across multiple executions.
In large model training, especially in heterogeneous computation, workers may not exhibit identical execution behavior due to differences in parallelization roles (e.g., pipeline/expert parallelism), input variability (e.g., variable-length video samples), and system effects (e.g., scheduling and contention). Nevertheless, worker-side execution is usually structured as repeated iterations that invoke a largely stable set of training-step functions and GPU kernels. Moreover, tensor shapes are often fixed by model configuration and batching policy, or are restricted to a small set of discrete shapes through padding, bucketing, truncation, or shape-constrained micro-batching. As a result, per-function runtime and resource-usage metrics across workers are expected to be either broadly consistent or, when not identical, to follow a relatively stable distribution reflecting systematic role differences and stochastic runtime noise. Under this assumption, function instances that fall outside the expected cross-worker distribution—i.e., statistically significant outliers in runtime or resource consumption—can be treated as anomalous behavior.